# GauRast: Enhancing GPU Triangle Rasterizers to Accelerate 3D Gaussian Splatting

### v1

@@ Abstract
3D intelligence leverages rich 3D features and stands as a promising frontier in AI, with 3D rendering fundamental to many downstream applications. 3D Gaussian Splatting (3DGS), an emerging high-quality 3D rendering method, requires significant computation, making real-time execution on existing GPU-equipped edge devices infeasible. Previous efforts to accelerate 3DGS rely on dedicated accelerators that require substantial integration overhead and hardware costs. This work proposes an acceleration strategy that leverages the similarities between the 3DGS pipeline and the highly optimized conventional graphics pipeline in modern GPUs. Instead of developing a dedicated accelerator, we enhance existing GPU rasterizer hardware to efficiently support 3DGS operations. Our results demonstrate a 23$\times$ increase in processing speed and a 24$\times$ reduction in energy consumption, with improvements yielding 6$\times$ faster end-to-end runtime for the original 3DGS algorithm and 4$\times$ for the latest efficiency-improved pipeline, achieving 24 FPS and 46 FPS respectively. These enhancements incur only a minimal area overhead of 0.2% relative to the entire SoC chip area, underscoring the practicality and efficiency of our approach for enabling 3DGS rendering on resource-constrained platforms.
@@
Hardware Architecture, Graphics Processors, 3D Gaussian Splatting
@@ I Introduction
Existing efforts to accelerate 3D Gaussian Splatting [16] and similar rendering pipelines [17, 10] typically focus on developing dedicated hardware accelerators. While these specialized units can provide performance benefits, they require exclusive hardware resources and often lead to increased system complexity and cost. Moreover, the integration of dedicated accelerators may not be feasible for all edge platforms, especially those with stringent area and power constraints. However, modern GPUs are already equipped with optimized fixed-function graphic acceleration hardware for triangle meshes. As these fixed-function units also target graphics rendering, their inherent capabilities present an opportunity to leverage existing resources to accelerate 3DGS rendering without introducing significant additional overhead.
@@
To mitigate the challenges associated with designing dedicated memory systems and software stacks for specialized accelerators [6], we propose an alternative solution: enhancing the existing graphics units—specifically, the rasterizer for triangle meshes—within GPUs to accelerate 3DGS rendering. By leveraging the capabilities of current GPU hardware, we aim to enhance performance without the need for additional dedicated accelerators. We begin with an in-depth profiling of the 3DGS rendering pipeline to identify time-consuming operations and analyze its similarities with operators in the triangle mesh rendering pipeline, which are already supported by the GPU’s native graphic hardware units. This analysis reveals opportunities to adapt and enhance these units to efficiently process 3D Gaussian operations.
@@
Building on these insights, we propose and develop an enhanced rasterizer, GauRast, for GPUs that enables the efficient execution of the dominant operator in the 3DGS rendering pipeline while preserving the original capabilities for standard triangle mesh rendering. This design ensures compatibility with existing GPU architectures and minimizes disruptions to conventional workflows. By enhancing existing hardware rather than introducing entirely new components, our approach strikes a balance between performance improvements and resource utilization. Our contributions are as follows:
@@
Step 2: Sorting. The second stage, as illustrated in Fig. 3(b), involves sorting the projected 2D Gaussians. Due to the potential overlap of Gaussians, each pixel may be influenced by multiple Gaussians. The rendering order impacts their visibility since Gaussians rendered earlier can obscure those rendered later. To maintain correct occlusion relationships and ensure visual consistency, the 2D Gaussians are sorted by their depth values. This depth-based sorting ensures that Gaussians nearer to the viewing position are rendered first, preserving proper occlusion handling.
Step 3: Gaussian Rasterization. The final step, Gaussian rasterization, applies the color of each Gaussian to the pixels it covers, based on opacity and Gaussian probability, and follows the depth order determined in Step 2. As shown in Fig. 3(c), for each Gaussian applied to a specific pixel, the density $\alpha_{P,i}$ is computed using the formula:
@@ III-A Opportunities in GPU’s Rasterizer
The computational demands of dominant Gaussian rasterization in the 3DGS rendering pipeline have motivated various acceleration efforts [16]. Current research primarily emphasizes the development of dedicated hardware accelerators for emerging 3D rendering pipelines such as 3DGS [16] and NeRF [17, 10]. Although these accelerators offer performance improvements, they can introduce significant trade-offs, including increased complexity in memory systems and software stacks [6] and the need for exclusive hardware resources. Unlike these dedicated approaches, our strategy focuses on enhancing existing GPU hardware to leverage the built-in capabilities of modern GPUs. This approach enhances 3DGS performance while preserving compatibility with conventional GPU operations and the original programming interface, minimizing disruptions to established workflows and fully utilizing the GPU’s existing data path.
Modern GPUs have dedicated units optimized for triangle mesh rasterization [23]. This process involves iterating over primitives (here referring to triangles) and mapping their parameters to display pixels. The pipeline is highly parallelized, relying on fixed-function accelerators for rapid computation [2, 5, 37]. Our approach begins with an investigation of the similarities and differences between Gaussian rasterization and triangle rasterization, the latter of which is efficiently supported by these optimized GPU hardware units.
@@ III-B Contrasting Gaussian and Triangle Rasterization Methods
However, there are also differences between the two rasterization processes. First, in triangle rasterization, the algorithm involves determining if a pixel lies within a triangular boundary. In contrast, 3DGS requires calculating Gaussian probability distributions to assess the coverage. This distinction necessitates modifications to the detection function to accommodate elliptical shapes rather than triangles. Second, for triangle mesh rasterization, the reduction method relies on a minimum-depth selection to determine which primitive’s parameters should be applied to each pixel, whereas 3D Gaussian Splatting aggregates color contributions from multiple overlapping Gaussians. This approach increases the complexity of the reduction function, as overlapping Gaussians contribute to pixel color based on their calculated densities and occlusion effects. Effectively addressing these differences is essential to adapting current rasterization hardware for 3DGS, ensuring compatibility and preserving performance.
@@ IV-B Proposed GauRast Hardware
Processing Element (PE). Each PE supports both Gaussian and triangle rasterization and is equipped with a combination of shared and dedicated logic units, as shown in Fig. 7(c). The shared logic handles common operations, while dedicated logic paths manage primitive-specific computations, such as evaluating Gaussian probability distributions for 3DGS rendering and performing depth division for triangle meshes. Importantly, the existing triangle rasterizer includes both shared and triangle-specific logic; thus, only Gaussian-specific logic is added. Compared to a standard triangle rasterizer, our design requires minimal additional hardware in each PE: two adders, one multiplier, and one exponentiation unit for Gaussian support, while reusing nine adders and nine multipliers from the triangle rasterizer. The PE is the only module requiring additional logic to facilitate 3DGS rendering.
@@ IV-C CUDA-Collaborative Scheduling
In triangle mesh rasterization, CUDA cores and the hardware rasterizer collaborate to complete the rendering process. Specifically, non-dominant tasks such as 3D-to-2D projection and color querying are handled by the CUDA cores, while the rasterization is managed by the hardware rasterizer [23]. Similarly, we adopt a hybrid scheduling strategy to enhance efficiency, assigning non-dominant operations like preprocessing and sorting to the CUDA cores, while the enhanced rasterizer handles the dominant rasterization workload which is the primary bottleneck in 3DGS rendering.
@@ V-A Evaluation Setup
V-B Comparison Against the Baseline SoC
Fig. 9 illustrates the speedup and efficiency improvements achieved by GauRast in Gaussian rasterization compared to CUDA implementations across all scenes in the benchmark dataset, for both the original algorithm [14] and its latest efficiency-optimized version [9]. The results show that our enhanced rasterizer achieves an average runtime reduction of 23$\times$ and an average energy efficiency improvement of 24$\times$ for the original algorithm. For the efficiency-optimized algorithm, it achieves a 20$\times$ runtime reduction and a 22$\times$ improvement in energy efficiency. Additionally, as shown in Fig. 10, replacing CUDA-based rasterization with our proposed enhanced rasterizer leads to substantial end-to-end performance improvements. GauRast achieves a 6$\times$ reduction in runtime for the original 3D Gaussian Splatting algorithm and a 4$\times$ reduction for the optimized algorithm, yielding average frame rates of 24 FPS and 46 FPS, respectively.
@@ V-C Comparison Against SOTA 3DGS Accelerator
[…]
@@ VII Acknowledgment
The authors would like to thank the anonymous reviewers for their valuable questions and constructive feedback. This work is supported by an internship at NVIDIA Research, the National Science Foundation (NSF) Computing and Communication Foundations (CCF) program (Award ID: 2312758), and the Department of Health and Human Services Advanced Research Projects Agency for Health (ARPA-H) under Award Number AY1AX000003.

### v2

@@ Abstract
3D intelligence leverages rich 3D features and stands as a promising frontier in AI, with 3D rendering fundamental to many downstream applications. 3D Gaussian Splatting (3DGS), an emerging high-quality 3D rendering method, requires significant computation, making real-time execution on existing GPU-equipped edge devices infeasible. Previous efforts to accelerate 3DGS rely on dedicated accelerators that require substantial integration overhead and hardware costs. This work proposes an acceleration strategy that leverages the similarities between the 3DGS pipeline and the highly optimized conventional graphics pipeline in modern GPUs. Instead of developing a dedicated accelerator, we enhance existing GPU rasterizer hardware to efficiently support 3DGS operations. Our results demonstrate a 23$\times$ increase in processing speed and a 24$\times$ reduction in energy consumption for the dominant rasterization operator in the 3DGS pipeline. These improvements yield a 6$\times$ and 4$\times$ faster end-to-end runtime for the original 3DGS algorithm and the latest efficiency-improved pipeline, respectively, achieving rendering speeds of 24 FPS and 46 FPS. These enhancements incur only a minimal area overhead of 0.2% relative to the entire SoC chip area, underscoring the practicality and efficiency of our approach for enabling 3DGS rendering on resource-constrained platforms.
@@
Architecture & System Design, Graphics Processors, 3D Gaussian Splatting, Neural Rendering
@@ I Introduction
Existing efforts to accelerate 3D Gaussian Splatting [17] and similar rendering pipelines [18, 11] typically focus on developing dedicated hardware accelerators. While these specialized units can provide performance benefits, they require exclusive hardware resources and often lead to increased system complexity and cost. Moreover, the integration of dedicated accelerators may not be feasible for all edge platforms, especially those with stringent area and power constraints. However, it is noteworthy that modern GPUs are already equipped with optimized fixed-function graphic acceleration hardware for triangle meshes. As these fixed-function units also target graphics rendering, their inherent capabilities present an opportunity to leverage existing resources to accelerate 3DGS rendering without introducing significant additional overhead.
@@
To mitigate the challenges associated with designing dedicated memory systems and software stacks for specialized accelerators [7], we propose an alternative solution: enhancing the existing graphics units—specifically, the rasterizer for triangle meshes—within GPUs to accelerate 3DGS rendering. By leveraging the capabilities of those existing fixed function units in current GPU hardware, we aim to enhance performance without the need for additional dedicated accelerators. We begin with an in-depth profiling of the 3DGS rendering pipeline to identify time-consuming operations and analyze its similarities with operators in the triangle mesh rendering pipeline, which are already supported by the GPU’s native graphic hardware units. This analysis reveals opportunities to adapt and enhance these units to efficiently process 3D Gaussian operations.
@@
Building on these insights, we propose and develop an enhanced rasterizer, GauRast, for GPUs that enables the efficient execution of the dominant operator in the 3DGS rendering pipeline while preserving the original capabilities for standard triangle mesh rendering. This design ensures compatibility with existing GPU architectures and minimizes disruptions to conventional workflows. By enhancing existing hardware rather than introducing entirely new components, our approach strikes a balance between performance improvements and resource overhead. Our contributions are as follows:
@@
Step 2: Sorting. The second stage, as illustrated in Fig. 3(b), involves sorting the projected 2D Gaussians. Due to the potential overlap of Gaussians, each pixel may be influenced by multiple Gaussians. The rendering order impacts their visibility since Gaussians rendered earlier can obscure those rendered later. To maintain correct occlusion relationships and ensure visual consistency, the 2D Gaussians are sorted by their depth values. This depth-based sorting ensures that Gaussians nearer to the viewing position are rendered first, preserving proper occlusion handling. Step 3: Gaussian Rasterization. The final step, Gaussian rasterization, applies the color of each Gaussian to the pixels it covers, based on opacity and Gaussian probability, and follows the depth order determined in Step 2. As shown in Fig. 3(c), for each Gaussian applied to a specific pixel, the density $\alpha_{P,i}$ is computed using the formula:
@@ III-A Opportunities in GPU’s Rasterizer
The computational demands of dominant Gaussian rasterization in the 3DGS rendering pipeline have motivated various acceleration efforts [17]. Current research primarily emphasizes the development of dedicated hardware accelerators for emerging 3D rendering pipelines such as 3DGS [17] and NeRF [18, 11]. Although these accelerators offer performance improvements, they can introduce significant trade-offs, including increased complexity in memory systems and software stacks [7] and the need for exclusive hardware resources. Unlike these dedicated approaches, our strategy focuses on enhancing existing GPU hardware to leverage the built-in capabilities of modern GPUs. This approach enhances 3DGS performance while preserving compatibility with conventional GPU operations and the original programming interface, minimizing disruptions to established workflows and fully utilizing the GPU’s existing data path. Modern GPUs have dedicated units optimized for triangle mesh rasterization [24]. This process involves iterating over primitives (here referring to triangles) and mapping their parameters to display pixels. The pipeline is highly parallelized, relying on fixed-function accelerators for rapid computation [2, 6, 39]. Our approach begins with an investigation of the similarities and differences between Gaussian rasterization and triangle rasterization, the latter of which is efficiently supported by these optimized GPU hardware units.
@@ III-B Contrasting Gaussian and Triangle Rasterization Methods
However, there are also differences between the two rasterization processes. First, in triangle rasterization, the algorithm involves determining if a pixel lies within a triangular boundary. In contrast, 3DGS requires calculating Gaussian probability distributions to assess the coverage. This distinction necessitates modifications to the detection function to accommodate elliptical shapes instead of being limited to triangles. Second, in triangle mesh rasterization, the reduction method relies on a simple minimum-depth comparison to determine the primitive with the minimum depth that should be applied to each pixel. In contrast, 3DGS aggregates color contributions from multiple overlapping Gaussians. This aggregation increases the complexity of the reduction function, as we need to compute densities and occlusion effects for all the overlapping Gaussians and then sum them together. Effectively addressing these differences is essential to adapting current rasterization hardware for 3DGS, ensuring compatibility and preserving performance.
@@ IV-B Proposed GauRast Hardware
Processing Element (PE). Each PE supports both Gaussian and triangle rasterization and is equipped with a combination of shared and dedicated logic units, as shown in Fig. 7(c). The shared logic handles common operations, while dedicated logic paths manage primitive-specific computations, such as evaluating Gaussian probability distributions for 3DGS rendering and performing depth division for triangle meshes. Importantly, the existing triangle rasterizer includes both shared and triangle-specific logic; thus, only Gaussian-specific logic is added. Compared to a standard triangle rasterizer, our design requires minimal additional hardware in each PE: two adders, one multiplier, and one exponentiation unit for Gaussian support, while reusing nine adders and nine multipliers from the triangle rasterizer. Multiplexers were employed to select the input for the hardware units based on the running mode. To save power, input gating is adopted when the modes do not match. The PE is the only module requiring additional logic to facilitate 3DGS rendering.
@@
Figure 8: Illustration of CUDA-Collaborative scheduling, where CUDA cores and the rasterizer operate in a pipelined manner to maximize hardware utilization during rendering.
@@ IV-C CUDA-Collaborative Scheduling
In triangle mesh rasterization, CUDA cores and the hardware rasterizer collaborate to complete the rendering process. Specifically, non-dominant tasks such as 3D-to-2D projection and color querying are handled by the CUDA cores, while the rasterization is managed by the hardware rasterizer [24]. Similarly, we adopt a hybrid scheduling strategy to enhance efficiency, assigning non-dominant operations like preprocessing and sorting to the CUDA cores, while the enhanced rasterizer handles the dominant rasterization workload which is the primary bottleneck in 3DGS rendering. Specifically, in the 3DGS’s 3-stage pipeline, the CUDA-cores process Stages 1-2 of one image before passing it to GauRast for Stage 3, after which the CUDA-cores can begin working on Stages 1-2 of the next image without waiting, as shown in Fig. 8.
@@ V-A Evaluation Setup
TABLE III: Absolute Rasterization Runtime w/ and w/o GauRast Compared to CUDA implementations on the NVIDIA Jetson Orin NX [23]. Evaluations were conducted using both the original 3DGS algorithm [15], across all seven scenes of the large-scale, real-world NeRF-360 dataset [4].
Scene Bicycle Stump Garden Room Counter Kitchen Bonsai
Baseline [23] 321 ms 149 ms 232 ms 236 ms 216 ms 269 ms 147 ms
GauRast 15 ms 6.0 ms 9.6 ms 10.5 ms 9.8 ms 12.2 ms 5.5 ms
@@
V-B Comparison Against the Baseline SoC
Fig. 10 and Tab. III illustrate the speedup and efficiency improvements achieved by GauRast in Gaussian rasterization compared to CUDA implementations across all scenes in the benchmark dataset, for both the original algorithm [15] and its latest efficiency-optimized version [10]. The results show that our enhanced rasterizer achieves an average runtime reduction of 23$\times$ and an average energy efficiency improvement of 24$\times$ for the original algorithm. For the efficiency-optimized algorithm, it achieves a 20$\times$ runtime reduction and a 22$\times$ improvement in energy efficiency. Additionally, as shown in Fig. 11, replacing CUDA-based rasterization with our proposed enhanced rasterizer leads to substantial end-to-end performance improvements. GauRast achieves a 6$\times$ reduction in runtime for the original 3D Gaussian Splatting algorithm and a 4$\times$ reduction for the optimized algorithm, yielding average frame rates of 24 FPS and 46 FPS, respectively.
@@ V-C Comparison Against SOTA 3DGS Accelerator
V-D Compatibility with non-NVIDIA GPUs
GauRast is compatible with any GPU equipped with a triangle rasterizer, including but not limited to those from NVIDIA. To further evaluate its generalizability, we conducted an experiment using an Apple M2 Pro GPU [3] with OpenSplat [31]. The M2 Pro provides 2.6× greater FP32 computing capability than the baseline NVIDIA Orin NX GPU [23]. Our results demonstrate that GauRast achieves an 11.2× speedup in rasterization on the bicycle scene, confirming its compatibility across different GPU architectures.
@@ VII Acknowledgment
The authors would like to thank Thierry Tambe and Rangharajan Venkatesan for their help with the HLS-based hardware design flow; Ben Eckart, Sifei Liu, and Marco Salvi for their participation in discussions on the 3DGS algorithm; and the anonymous reviewers for their valuable questions and constructive feedback. This work is supported by an internship at NVIDIA Research, the National Science Foundation (NSF) Computing and Communication Foundations (CCF) program (Award ID: 2312758), and the Department of Health and Human Services Advanced Research Projects Agency for Health (ARPA-H) under Award Number AY1AX000003.